# Sharp Estimates for the Robin Laplacian under a Perimeter Constraint in hyperbolic space

### v1

@@ Abstract.
In this paper, we establish a lower bound, in terms of the isoperimetric deficit, for the first eigenvalue of the Robin Laplacian with negative boundary parameter on horospherically convex bounded domains in hyperbolic space, which implies that the geodesic ball maximizes this eigenvalue among all such domains. Furthermore, we derive upper bounds for the first eigenvalue of the Robin Laplacian with positive boundary parameter on horospherically convex bounded domains in hyperbolic space.
@@ 1. Introduction
In [ACC+25], Acampora et al. established a lower bound for the first eigenvalue of the Robin Laplacian with negative parameter on strongly convex sets in the sphere $\mathbb{S}^{n}$, under a perimeter constraint. The authors [ACC+25] also constructed convex sets for which the inner parallel sets are not convex in hyperbolic space $\mathbb{H}^{n}$. Motivated by [ACC+25], we observe that the inner parallel sets of horospherically convex bounded domains are also horospherically convex in $\mathbb{H}^{n}$. Based on this observation, in the first part, for $\beta<0$, we establish a lower bound for $\lambda_{1}\left(\Omega,\beta\right)$ on horospherically convex bounded domains in the hyperbolic space $\mathbb{H}^{n}$ by means of the isoperimetric deficit.
@@
According to the isoperimetric inequality in hyperbolic space [Cha84], the following result holds.
@@
The paper is organized as follows. In Section 2, we recall some notions, the Steiner formula and the Alexandrov-Fenchel inequality in hyperbolic space. This section also analyzes the convexity of the parallel domains and provides an estimate for the perimeter of the inner parallel domains. In the final section, we present complete proofs of Theorems 1.1 and 1.2.
@@ 2. Preliminaries
This section recalls some basic notions and properties in hyperbolic space. For details, we refer to [Koh91, WX14]. Furthermore, we analyze the convexity of parallel domains and derive an estimate for the perimeter of the inner parallel domains.

### v2

@@ Abstract.
In this paper, we establish a lower bound, in terms of the isoperimetric deficit, for the first eigenvalue of the Robin Laplacian with negative boundary parameter on horospherically convex bounded domains in the hyperbolic space. This implies that the geodesic ball maximizes this eigenvalue among all such domains, thereby providing a partial resolution to an open problem posed by Celentano, Krejčiřík and Lotoreichik in [CKL26]. Furthermore, we derive upper bounds for the first eigenvalue of the Robin Laplacian with positive boundary parameter on horospherically convex bounded domains in the hyperbolic space.
@@ 1. Introduction
In [ACC+25], Acampora et al. established a lower bound for the first eigenvalue of the Robin Laplacian with negative parameter on strongly convex sets in the sphere $\mathbb{S}^{n}$, under a perimeter constraint. The authors [ACC+25] also constructed convex sets for which the inner parallel sets are not convex in the hyperbolic space $\mathbb{H}^{n}$. In the introduction of [CKL26], Celentano, Krejčiřík and Lotoreichik stated $\cdots$, to the best of our knowledge, it remains an open problem whether among all bounded smooth simply-connected domains in the hyperbolic plane with fixed perimeter the lowest eigenvalue of the Robin Laplacian with negative boundary parameter is maximised by the geodesic disk.” We observe that the inner parallel sets of horospherically convex bounded domains are also horospherically convex in $\mathbb{H}^{n}$. Based on this observation, in the first part, for $\beta<0$, we establish a lower bound for $\lambda_{1}\left(\Omega,\beta\right)$ on horospherically convex bounded domains in hyperbolic space $\mathbb{H}^{n}$ by means of the isoperimetric deficit. This bound implies that the geodesic ball maximizes this eigenvalue, thereby providing a partially affirmative answer to the open problem posed in [CKL26].
@@
According to the isoperimetric inequality in the hyperbolic space [Cha84], the following result holds.
@@
The paper is organized as follows. In Section 2, we recall some notions, the Steiner formula and the Alexandrov-Fenchel inequality in the hyperbolic space. This section also analyzes the convexity of the parallel domains and provides an estimate for the perimeter of the inner parallel domains. In the final section, we present complete proofs of Theorems 1.1 and 1.2.
@@ 2. Preliminaries
This section recalls some basic notions and properties in the hyperbolic space. For details, we refer to [Koh91, GWW14, WX14]. Furthermore, we analyze the convexity of parallel domains and derive an estimate for the perimeter of the inner parallel domains.
@@ 2.B. Parallel hypersurfaces and Alexandrov-Fenchel inequality
Acknowledgements
The authors thank Professor Yong Wei for his helpful discussions.